# Phase separation of chemokinetic active particles

### v1

@@ IV.2 Quorum-sensing ABPs
By evaluating this quantity, we can determine which hydrodynamic model corresponds to a given system of quorum-sensing ABPs. The second advantage is that, since there is no short-range repulsion, the simulation of quorum-sensing ABPs can work even with relatively large time steps without blowing up. Indeed, we use $\Delta t=10^{-3}$ for the quorum-sensing ABPs, while $\Delta t=10^{-5}$ has to be used for the mechanically interacting ABPs.
Figure 9: Dependence of the time-variability $\mathcal{V}(\tau)$ on $D_{\mathrm{c}}$ in the AMR model of QS simulation, evaluated with $\tau=200$, $L_{\mathrm{win}}=35$. Other parameters: $\lambda=0.0025$, $n_{0}=20$, $\zeta=22$. Inset: mean (round marker) and standard deviation (error bar) of cluster sizes for $n_{0}=20$, $\zeta=22$, $D_{\mathrm{c}}=2.5$ at $t=2\times 10^{3}$.
Let us now examine how the behaviors of the quorum-sensing ABPs compare with the hydrodynamic model. For the BMR, as detailed in Appendix A, the boundary between the H and S phases is determined only by infinite-wavelength features of the system. The property, which is also reflected in the absence of $\kappa$ in Eq. (10) for the phase boundary of the BMR, implies that the phase boundary of the quorum-sensing ABPs coincides exactly with that of the hydrodynamic theory. This is verified by Fig. 8(a), which shows the snapshots of the simulations of the quorum-sensing ABPs at $t=4\times 10^{2}$ (starting from uniform configurations at $t=0$) together with the phase boundary predicted by Eq. (10) (marked by the dotted line). As was the case for the hydrodynamic model (see Fig. 2), the system tends to form more concentrated particle clusters for higher $n_{0}$, while sparse regions broaden at the expense of dense regions as $n_{0}$ is decreased. If the system is allowed to evolve even longer (to $t=4\times 10^{2}$), different clusters merge via the Ostwald process, forming large-scale patterns (see Fig. S2) also observed in the hydrodynamic model.
For the AMR, the short-wavelength features do affect the phase boundaries, so the correspondence between the quorum-sensing ABPs and the hydrodynamic model cannot be exact. Still, Fig. 8(b) and Supplementary Movies XXX suggest some qualitative agreements between the two. At fixed $D_{\mathrm{c}}$, the snapshots show that the particle clusters shrink with increasing $\lambda$. This is also confirmed by the $\lambda$-dependence of the mean cluster size at $D_{\mathrm{c}}=2.5$ shown in the inset of Fig. 9, which also reveals that the mean cluster size does not change appreciably with the system size $L$. These results are consistent with the existence of a transition from the S phase with microphase separation to the H phase predicted by the hydrodynamic theory.
@@
where $\bar{n}(\mathbf{r},t)$ is the average of the chemical concentration $n(\mathbf{r},t)$ over a square of side length $L_{\mathrm{win}}$ centered at the position $\mathbf{r}$. By examining how large $\mathcal{V}_{\tau}$ can be for a suitable choice of $\tau$ and $L_{\mathrm{win}}$, one can quantify the magnitude, time scale, and length scale of oscillations present in the system. At $\lambda=0.0025$, choosing $\tau=200$ and $L_{\mathrm{win}}=35$, we observe a dramatic increase of $\mathcal{V}_{\tau}$ as $D_{\mathrm{c}}$ is lowered [see Fig. 9]. Moreover, the behaviors of $\mathcal{V}_{\tau}$ also suggest that the length scale of cluster patterns tends to decrease with increasing $\lambda$, as shown in Fig. S3. These support the existence of the S–O and the H–O boundaries predicted by the hydrodynamic theory.
@@
We note that the ballistic bands and the radiating ripples generated by the hydrodynamic model are not clearly observable for the quorum-sensing ABPs. This may be due to the microscopic fluctuations, which are not fully removed by introducing mean-field interactions, destroying such large-scale patterns. Still, our results thus far imply that a particle-based model can exhibit all of the three phases predicted by the hydrodynamic theory.
V Summary and outlook
We showed that chemokinesis combined with chemical consumption can profoundly affect the collective behaviors of active particles. For the BMR, the chemical consumption is increased inside dense particle clusters, resulting in lower chemical concentration there and more robust clusters compared to the vanilla case. Conversely, when particles in the AMR are clustered, the enhanced chemical consumption within the region reduces cluster size or even leads to the evaporation of temporary clusters. We validated our scenarios through three approaches: linear stability analysis, numerical simulations of the coarse-grained hydrodynamic equations, and particle simulations.
@@ V Summary and outlook
To our best knowledge, this study is the first to demonstrate the importance of chemical consumption mechanism on the patterns of active matter. Future researches may contain similar analyses for models beyond scalar active matter, including models with polar order, topological defects, etc. Moreover, various mechanical abilities of active matter, such as enhanced transport and current rectification may also be modified by chemical consumption. This may affect the performance of collective machines utilizing active matter, and therefore should be investigated further.
@@
Acknowledgments. — This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00218318).
Appendix A Linear stability analysis
@@
Here we provide details of the LSA whose results are used in Sec. III. Suppose that the density $\rho$ and the concentration $n$ are weakly perturbed around their homogeneous values $\rho_{0}$ and $n_{0}$, respectively:
$\rho=\rho_{0}+\delta\rho,\,\,n=n_{0}+\delta n\;.$ (S1)
@@
Using these in Eq. (II), the time evolution of $\delta\rho$ is expanded to linear order as
@@
$\delta\dot{\rho}=D_{\text{eff}}\nabla^{2}\delta\rho-\kappa\nabla^{2}(\nabla^{2}\delta\rho)+\frac{\alpha\rho_{0}(\alpha n_{0}-\rho_{0}\zeta)}{2D_{\mathrm{r}}}\nabla^{2}\delta n\;,$ (S2)
@@
$D_{\text{eff}}\equiv D+\frac{(\alpha n_{0}-\rho_{0}\zeta)(\alpha n_{0}-2\rho_{0}\zeta)}{2D_{\mathrm{r}}}$ (S3)
@@
$\delta\dot{n}=D_{\mathrm{c}}\nabla^{2}\delta n-\lambda n_{0}\,\delta\rho-\lambda\rho_{0}\,\delta n$ (S4)
@@
$\delta\dot{n}$ $=D_{\mathrm{c}}\nabla^{2}\delta n-\lambda n_{0}(\alpha n_{0}-2\rho_{0}\zeta)\,\delta\rho$
$\qquad-\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)\,\delta n$ (S5)
@@
Now, taking the Fourier components $\delta\rho=\delta\rho_{0}e^{i\mathbf{q}\cdot\mathbf{r}+\omega t}$ and $\delta n=\delta n_{0}e^{i\mathbf{q}\cdot\mathbf{r}+\omega t}$, Eqs. (S2), (S4), and (A) can be rewritten as an eigenvalue equation
$\begin{pmatrix}\mathcal{A}&\mathcal{B}\\ \mathcal{C}&\mathcal{D}\end{pmatrix}\begin{pmatrix}\delta\rho_{0}\\ \delta n_{0}\end{pmatrix}=\omega\begin{pmatrix}\delta\rho_{0}\\ \delta n_{0}\end{pmatrix}\;,$ (S6)
@@
$\mathcal{D}$ =$=\begin{cases}-D_{\mathrm{c}}q^{2}-\lambda\rho_{0}&\text{(BMR)}\\ -D_{\mathrm{c}}q^{2}-\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)&\text{(AMR)}\end{cases}\;.$ (S7)
@@
$\omega^{2}-(\mathcal{A}+\mathcal{D})\omega+\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}=0\;,$ (S8)
@@
$\omega_{\pm}=\frac{1}{2}\left[(\mathcal{A}+\mathcal{D})\pm\sqrt{(\mathcal{A}-\mathcal{D})^{2}+4\mathcal{B}\mathcal{C}}\right]\;.$ (S9)
@@
A.1 Case of the BMR
In this case, $\mathcal{B}\mathcal{C}$ is always positive. Then Eq. (S9) ensures that both eigenvalues are real, with the boundary between the H and the S phases determined by when $\omega_{+}$ changes its sign. In the S phase, near the phase boundary, we must have $\omega_{+}>0$ and $\omega_{-}<0$. By examining the quadratic left-hand side of Eq. (S8), one can see that this is equivalent to the condition $\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}<0$. By simple algebra, we obtain $\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}=q^{2}h(q^{2})$, where
$h(q^{2})$ $=\kappa D_{\mathrm{c}}q^{4}+(D_{\mathrm{eff}}D_{\mathrm{c}}+\kappa\lambda\rho_{0})q^{2}+\lambda\rho_{0}D_{\mathrm{eff}}$
$\quad-\frac{\lambda\rho_{0}}{2D_{\mathrm{r}}}\alpha n_{0}(\alpha n_{0}-\rho_{0}\zeta)$ (S10)

### v2

@@ Results
For a “fair” comparison with the original hydrodynamic model, we can set $\kappa$ of the model equal to the value determined by the above equation. The second advantage is that, since there is no short-range repulsion, the simulation of quorum-sensing ABPs can work even with relatively large time steps without blowing up. Indeed, we use $\Delta t=10^{-3}$ for the quorum-sensing ABPs, while $\Delta t=10^{-5}$ has to be used for the mechanically interacting ABPs.
Figure 7: Properties of patterns formed by the quorum-sensing active Brownian particles in the Active Metabolic Regime. The main figure shows how the time-variability $\mathcal{V}_{\tau}$ depends on the chemical diffusion constant $D_{\mathrm{c}}$ for $\tau=200$, $L_{\mathrm{win}}=35$, and $\lambda=0.0025$. Inset: The mean (symbols) and the standard deviation (error bars) of cluster sizes as functions of the chemical consumption rate $\lambda$ and the system size $L$ for $D_{\mathrm{c}}=2.5$ at $t=2\times 10^{3}$. For all plots, we use $n_{0}=20$ and $\zeta=22$.
Let us now examine how the behaviors of the quorum-sensing ABPs compare with the hydrodynamic model. For the BMR, as detailed in Methods, the boundary between the H and S phases is determined only by infinite-wavelength features of the system. The property, which is also reflected in the absence of $\kappa$ in Eq. (11) for the phase boundary of the BMR, implies that the phase boundary of the quorum-sensing ABPs coincides exactly with that of the hydrodynamic theory. This is verified by Fig. 6(a), which shows the snapshots of the simulations of the quorum-sensing ABPs at $t=4\times 10^{2}$ (starting from uniform configurations at $t=0$) together with the phase boundary predicted by Eq. (11) (marked by the dotted line). As was the case for the hydrodynamic model (see Fig. A4), the system tends to form more concentrated particle clusters for higher $n_{0}$, while sparse regions broaden at the expense of dense regions as $n_{0}$ is decreased. If the system is allowed to evolve even longer (to $t=2\times 10^{3}$), different clusters merge via the Ostwald process, forming large-scale patterns (see Fig. A5 and Appendix E) also observed in the hydrodynamic model.
For the AMR, the short-wavelength features (which breaks the assumption of slow spatial variations used in the hydrodynamic theory) affect the phase boundaries, so the correspondence between the quorum-sensing ABPs and the hydrodynamic model cannot be exact. Still, Fig. 6(b) and Supplementary Movie 4B suggest some qualitative agreements between the two. At fixed $D_{\mathrm{c}}$, the snapshots show that the particle clusters shrink with increasing $\lambda$. This is also confirmed by the $\lambda$-dependence of the mean cluster size at $D_{\mathrm{c}}=2.5$ shown in the inset of Fig. 7, which also reveals that the mean cluster size does not change appreciably with the system size $L$. These results are qualitatively similar to how the S phase changes to the H phase in the hydrodynamic theory. The reason why the H phase is not observed within the simulated range of $\lambda$ may be an artifact generated by Eq. (17), which completely freezes small high-density regions satisfying $\alpha n-\varrho_{\mathrm{loc}}\zeta<0$.
@@
where $\bar{n}(\mathbf{r},t)$ is the average of the chemical concentration $n(\mathbf{r},t)$ over a square window of side length $L_{\mathrm{win}}$ centered at the position $\mathbf{r}$. By examining how large $\mathcal{V}_{\tau}$ can be for a suitable choice of $\tau$ and $L_{\mathrm{win}}$, one can quantify the magnitude, time scale, and length scale of cluster fluctuations present in the system. At $\lambda=0.0025$, choosing $\tau=200$ and $L_{\mathrm{win}}=35$, we observe a dramatic increase of $\mathcal{V}_{\tau}$ as $D_{\mathrm{c}}$ is lowered [see Fig. 7]. While $\mathcal{V}_{\tau}$ does not exhibit the dip near the S–O boundary shown by the mean vorticity in Fig. 4(f), this may be because the quantity is not as sensitive to the total length of cluster boundaries as the mean vorticity is. Moreover, the behaviors of $\mathcal{V}_{\tau}$ also suggest that the length scale of cluster patterns tends to decrease with increasing $\lambda$, as shown in Fig. A6 and Appendix E. These support the existence of the S–O and the H–O boundaries predicted by the hydrodynamic theory.
@@
We note that the ballistic bands and the radiating ripples generated by the hydrodynamic model are not observable for the quorum-sensing ABPs. It might be that such large-scale patterns are destroyed by the higher-order gradients neglected in the hydrodynamic theory or the microscopic fluctuations which are not entirely removed by introducing mean-field interactions. Still, our results thus far imply that particle-based models can exhibit both S and O phases predicted by the hydrodynamic theory.
Discussion
We showed that chemokinesis combined with chemical consumption can profoundly affect the collective behaviors of active particles. In the BMR, where only a negligible portion of chemical consumption is related to particle motion, lower chemical concentration and weaker self-propulsion within particle clusters strengthen the tendency for MIPS. In the AMR, where chemical consumption is tightly coupled to particle motion, MIPS is suppressed due to higher chemical concentration and stronger self-propulsion within particle clusters. In this regime, MIPS occurs only as microphase separation. According to the hydrodynamic theory, this regime is divided further into the S phase, where the particle clusters are stable, and the O phase, where the clusters form, move, and vanish rapidly. While we are not aware of any experimental studies addressing the interplay of MIPS, chemokinesis, and chemical consumption, the predictions of our theory would be relevant to synthetic active particles with predominantly chemokinetic response to the fuel [45, 46].
We also note that the AMR’s O phase resembles the dynamic clustering observed in the colloidal suspension of platinum–gold Janus particles [31], which was attributed to their chemotactic response to the fuel [28, 29, 30]. However, the O phase requires that the relaxation time scale of chemical concentration is comparable to or slower than that of particle density. To meet this criterion, the diffusion constant $D_{\mathrm{c}}$ of the chemical must be less than or comparable to the effective diffusion constant $D_{\mathrm{eff}}$ of the active particles. Experimental realizations (see [3] and the references therein) of micron-size Janus particles in an aqueous solution of the fuel (e.g., hydrogen peroxide) typically satisfy $D_{\mathrm{eff}}\approx 1~{}\mu\mathrm{m}^{2}s^{-1}$ and $D_{\mathrm{c}}\approx 10^{3}~{}\mu\mathrm{m}^{2}s^{-1}$, so the O phase is well beyond the regime of such empirical systems. To achieve $D_{\mathrm{c}}\lesssim D_{\mathrm{eff}}$, we must either reduce the size of the active particle to nanoscale (active particles as small as $50~{}\mathrm{nm}$ were achieved in [67, *Ma2015enzyme]) or increase the self-propulsion velocity $v$ by a factor of $30$ (since $D_{\mathrm{eff}}\sim v^{2}$).
The dichotomy between the full phase separation of the BMR and the microphase separation of the AMR is reminiscent of the same two types of phase separations exhibited by the Active Model B+ [16]. Whether there exists a common physical principle underlying such similar dichotomies is an interesting question. Moreover, it remains to be checked whether some higher-order terms or microscopic fluctuations neglected in the hydrodynamic theory destroy the large-scale oscillating patterns predicted for the AMR.
@@ Discussion
This study demonstrates the importance of chemical consumption mechanisms on the patterns of active matter. Future research may contain similar analyses for models beyond scalar active matter, including models with polar order, topological defects, etc. Moreover, chemical consumption may also modify various mechanical abilities of active matter, such as enhanced transport and current rectification. This may affect the performance of collective machines utilizing active matter and, therefore, should be investigated further.
Data availability
The data that support the plots within this paper and other findings of this study are available from the corresponding author upon reasonable request.
Code availability
The computer codes used to generate the results of this paper are available at https://github.com/KwonEuiJoon/chemokinetic_active_matter.
Acknowledgments
@@
This work was supported by the National Research Foundation of Korea (NRF) grants (RS-2021-017476, RS-2023-00218318, and RS-2023-00278985, RS-2024-00410147) funded by the Ministry of Science and ICT (MSIT) of the Korea government. We thank Étienne Fodor and Ignacio Pagonabarraga for helpful discussions.
Author contributions
E.K. and Y.O. derived and analyzed the hydrodynamic equations and performed numerical simulations. Y.B. supervised the project. All authors participated in the writing of the manuscript.
Competing interests
The authors declare no competing interests.
Appendix A Derivation of hydrodynamic equations
Here we provide the details of how the hydrodynamic equations [Eq. (Hydrodynamic model) in the main text] are derived from the particle-based models [Eqs. (Particle-based model4) in the main text]. The procedure closely follows the coarse-graining scheme given in [12]. Denoting by $\varphi_{k}$ the orientational angle of the self-propulsion direction $\mathbf{e}_{k}=(\cos\varphi_{k},\sin\varphi_{k})$ of particle $k$, the time evolution of the normalized joint probability distribution $\psi_{N}(\{\mathbf{r}_{k},\varphi_{k}\};t)$ is governed by
$\partial_{t}\psi_{N}=\sum_{k=1}^{N}\nabla_{k}\cdot[\mu(\nabla_{k}U)-\tilde{v}\mathbf{e}_{k}+D_{0}\nabla_{k}]\psi_{N}+D_{\mathrm{r}}\sum_{k=1}^{N}\frac{\partial^{2}\psi_{N}}{\partial\varphi_{k}^{2}}\;,$ (A1)
where $D_{0}=\mu T$ and $U=\sum_{k\neq k^{\prime}}V(|\mathbf{r}_{k}-\mathbf{r}_{k^{\prime}}|)$. The first step is to integrate out the degrees of freedom associated with $N-1$ particles, which yields the projected density
[…]
of a single particle. Since the particles are identical, from now on we drop the subscripts of $\mathbf{r}_{1}$ and $\varphi_{1}$, using $\mathbf{r}$ and $\varphi$ instead. Integrating Eq. (A1) side by side, we get
$\partial_{t}\psi_{1}=-\nabla\cdot[\mu\mathbf{F}+\tilde{v}\,\mathbf{e}\,\psi_{1}-D_{0}\nabla\psi_{1}]+D_{\mathrm{r}}\partial^{2}_{\varphi}\psi_{1}\;,$ (A3)
with $\mathbf{e}\equiv(\cos\varphi,\sin\varphi)$ and the mean interaction force given by
[…]
Here, $\psi_{2}(\mathbf{r},\varphi,\mathbf{r}^{\prime};t)$ is the two-body density associated with the joint probability that there is a particle of orientation $\varphi$ at $\mathbf{r}$ while another is located at $\mathbf{r}^{\prime}$.
Next, we aim to formally eliminate the dependence of Eq. (A3) on $\mathbf{F}$, replacing it with the dependence on the local particle density
[…]
Then, the two-particle distribution can be reexpressed as
$\psi_{2}(\mathbf{r},\varphi,\mathbf{r}^{\prime}\,;\,t)=\psi_{1}(\mathbf{r},\varphi\,;\,t)\,\rho(\mathbf{r})\,g(\mathbf{r}^{\prime}\,|\,\mathbf{r},\varphi\,;\,t)\;,$ (A6)
where $g(\mathbf{r}^{\prime}\,|\,\mathbf{r},\varphi\,;\,t)$ is the conditional probability of finding another particle at $\mathbf{r}^{\prime}$ given that there is a particle at $\mathbf{r}$ with orientation $\varphi$. Further simplifications can be made by assuming that the system is effectively homogeneous, isotropic, and fast-relaxing when it comes to the two-particle statistics (despite the spatial and temporal variations of $\rho$ and $\mathbf{p}$). Under these assumptions, we can write
$g(\mathbf{r}^{\prime}\,|\,\mathbf{r},\varphi\,;\,t)\simeq g(|\mathbf{r}-\mathbf{r}^{\prime}|,\theta)\;,$ (A7)
so that only the distance $|\mathbf{r}-\mathbf{r}^{\prime}|$ and the directional angle $\theta$ matter. Then, we get $\mathbf{e}\cdot\mu\mathbf{F}=-\rho\zeta\psi_{1}$ with
[…]
We may assume that $\phi_{1}$ is slow-varying in space, which allows us to decompose $\mathbf{F}$ linearly in terms of the orientation $\mathbf{e}$ and the density gradient $\nabla\psi_{1}$. This yields
$\mathbf{F}=(\mathbf{e}\cdot\mathbf{F})\mathbf{e}+(D_{0}-D)\nabla\psi_{1}+\mathcal{O}(|\nabla\psi_{1}|^{2})\;,$ (A9)
where $D\equiv D_{0}-\mathbf{F}\cdot[\nabla\psi_{1}-\mathbf{e}(\mathbf{e}\cdot\nabla\psi_{1})]/|\nabla\psi_{1}|$. Using the relation in Eq. (A3), the equation of motion for $\psi_{1}$ is simplified to
$\partial_{t}\psi_{1}=-\nabla\cdot[v\mathbf{e}-D\nabla]\psi_{1}+D_{\mathrm{r}}\partial^{2}_{\varphi}\psi_{1}\;,$ (A10)
where $v\equiv\tilde{v}-\rho\zeta$. Thus, we have formally replaced the effects of the interaction force $\mathbf{F}$ with the deceleration effect governed by $-\rho\zeta$ and the renormalized diffusion coefficient $D$. While $D$ defined above is in principle not uniform in space, we resort to the effective homogeneity of the system and treat $D$ as a constant in the hydrodynamic equations.
While Eq. (A10) has no explicit dependence on $\mathbf{F}$, it is still dependent on the particle orientation $\mathbf{e}$. To eliminate the dependence, we define the polarization field
[…]
Then, integrating Eq. (A10) side by side with respect to $\varphi$ before and after multiplying $\mathbf{e}$, we obtain
$\partial_{t}\rho(\mathbf{r},t)$ $=-\nabla\cdot[v\mathbf{p}(\mathbf{r},t)-D\nabla\rho(\mathbf{r},t)]$ (A12)
$\partial_{t}\mathbf{p}(\mathbf{r},t)$ $=-\frac{1}{2}\nabla(v\rho)+D\nabla^{2}\mathbf{p}-D_{\mathrm{r}}\mathbf{p}\;,$ (A13)
where we neglected the nematic contributions assuming that they quickly relax to zero. We further assume that $\mathbf{p}(\mathbf{r},t)$ also quickly relaxes to the steady state determined by $\rho(\mathbf{r},t)$, with slow spatial variations. Then, at the leading order, we can use the approximation $\mathbf{p}(\mathbf{r},t)\simeq-\nabla(v\rho)/(2D_{\mathrm{r}})$. Plugging this expression into Eq. (A12) yields
$\partial_{t}\rho=\nabla\cdot\left[D\nabla\rho+\frac{v}{2D_{\mathrm{r}}}\nabla(v\rho)\right]\;.$ (A14)
ion To obtain Eq. (Hydrodynamic model) in the main text, we need to introduce an additional term $\kappa\nabla^{2}(\nabla^{2}\rho)$ on the right-hand side of the above equation, which is the next lowest-order term allowed by the symmetry of the system and phenomenologically necessary to account for the interface energy between dense and dilute regions. Microscopically, the term would naturally arise from higher-order gradients of the fields that were neglected in this derivation.
Now, we derive the hydrodynamic equation for the chemical. Assuming that sufficiently many active Brownian particles (ABPs) interact with the chemical at every location, the chemical concentration $n(\mathbf{r},t)$ can be regarded as a deterministic field, with the consumption function $f$ in Eq. (2) in the main text replaced with its average over the particle distribution $\psi_{N}$, which we denote by $\langle f\rangle$. For the Basal Metabolic Regime (BMR), it is straightforward to see that
$\langle f(n,\{\mathbf{r}_{k},\dot{\mathbf{r}}_{k}\})\rangle=n(\mathbf{r})\sum_{k^{\prime}=1}^{N}\langle\delta(\mathbf{r}-\mathbf{r}_{k^{\prime}})\rangle=n(\mathbf{r})\rho(\mathbf{r})\;.$ (A15)
For the Active Metabolic Regime (AMR), we rewrite $\dot{\mathbf{r}}_{k}$ in terms of the three terms on the right-hand side of Eq. (Particle-based model) of the main text and calculate the contribution of each term to $\langle f\rangle$ separately. The first term contributes
$\sum_{k^{\prime}=1}^{N}\left\langle\tilde{v}(n(\mathbf{r}_{k^{\prime}})))\delta(\mathbf{r}-\mathbf{r}_{k^{\prime}})\right\rangle=\tilde{v}(n(\mathbf{r}))\rho(\mathbf{r})\;.$ (A16)
Meanwhile, the second term contributes
$\sum_{k\neq k^{\prime}}\mu\left\langle\mathbf{e}_{k}\cdot\nabla_{k}V(|\mathbf{r}_{k}-\mathbf{r}_{k^{\prime}}|)\delta(\mathbf{r}-\mathbf{r}_{k})\right\rangle$
[…]
$=\rho(\mathbf{r})^{2}\zeta\;,$ (A17)
where we used the relation above Eq. (A8) to obtain the third equality. Finally, the third term has zero contribution as the white noise $\bm{\xi}_{k}$ averages to zero. Combining the above results, the AMR satisfies
$\langle f(n,\{\mathbf{r}_{k},\dot{\mathbf{r}}_{k}\})\rangle=n\rho(\tilde{v}-\rho\zeta)=n\rho v(\rho,n)\;.$ (A18)
Using Eqs. (3) and (4) to approximate Eq. (2) in the main text, we obtain the hydrodynamic equation for the chemical concentration $n(\mathbf{r},t)$ shown in Eq. (Hydrodynamic model).
Appendix B Details of linear stability analysis
@@
Here we provide details of linear stability analysis (LSA) for the prediction of MIPS. Suppose that the density $\rho$ and the concentration $n$ are weakly perturbed around their homogeneous values $\rho_{0}$ and $n_{0}$, respectively:
$\rho=\rho_{0}+\delta\rho,\,\,n=n_{0}+\delta n\;.$ (A19)
@@
Using these in Eq. (7) of the main text, the time evolution of $\delta\rho$ is expanded to linear order as
@@
$\delta\dot{\rho}=D_{\text{eff}}\nabla^{2}\delta\rho-\kappa\nabla^{2}(\nabla^{2}\delta\rho)+\frac{\alpha\rho_{0}(\alpha n_{0}-\rho_{0}\zeta)}{2D_{\mathrm{r}}}\nabla^{2}\delta n\;,$ (A20)
@@
$D_{\text{eff}}\equiv D+\frac{(\alpha n_{0}-\rho_{0}\zeta)(\alpha n_{0}-2\rho_{0}\zeta)}{2D_{\mathrm{r}}}$ (A21)
@@
$\delta\dot{n}=D_{\mathrm{c}}\nabla^{2}\delta n-\lambda n_{0}\,\delta\rho-\lambda\rho_{0}\,\delta n$ (A22)
@@
$\delta\dot{n}=D_{\mathrm{c}}\nabla^{2}\delta n-\lambda n_{0}(\alpha n_{0}-2\rho_{0}\zeta)\,\delta\rho-\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)\,\delta n$ (A23)
@@
Now, taking the Fourier components $\delta\rho=\delta\rho_{0}e^{i\mathbf{q}\cdot\mathbf{r}+\omega t}$ and $\delta n=\delta n_{0}e^{i\mathbf{q}\cdot\mathbf{r}+\omega t}$, Eqs. (A20), (A22), and (A23) can be rewritten as an eigenvalue equation
$\begin{pmatrix}\mathcal{A}&\mathcal{B}\\ \mathcal{C}&\mathcal{D}\end{pmatrix}\begin{pmatrix}\delta\rho_{0}\\ \delta n_{0}\end{pmatrix}=\omega\begin{pmatrix}\delta\rho_{0}\\ \delta n_{0}\end{pmatrix}\;,$ (A24)
@@
$\mathcal{D}$ =$=\begin{cases}-D_{\mathrm{c}}q^{2}-\lambda\rho_{0}&\text{(BMR)}\\ -D_{\mathrm{c}}q^{2}-\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)&\text{(AMR)}\end{cases}\;.$ (A25)
@@
$\omega^{2}-(\mathcal{A}+\mathcal{D})\omega+\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}=0\;,$ (A26)
@@
$\omega_{\pm}=\frac{1}{2}\left[(\mathcal{A}+\mathcal{D})\pm\sqrt{(\mathcal{A}-\mathcal{D})^{2}+4\mathcal{B}\mathcal{C}}\right]\;.$ (A27)
@@
B.0.1 Case of the BMR
In this case, $\mathcal{B}\mathcal{C}$ is always positive. Then Eq. (A27) ensures that both eigenvalues are real, with the boundary between the H and the S phases determined by when $\omega_{+}$ changes its sign. In the S phase, near the phase boundary, we must have $\omega_{+}>0$ and $\omega_{-}<0$. By examining the quadratic left-hand side of Eq. (A26), one can see that this is equivalent to the condition $\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}<0$. By simple algebra, we obtain $\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}=q^{2}h(q^{2})$, where
$h(q^{2})=\kappa D_{\mathrm{c}}q^{4}+(D_{\mathrm{eff}}D_{\mathrm{c}}+\kappa\lambda\rho_{0})q^{2}+\lambda\rho_{0}D_{\mathrm{eff}}-\frac{\lambda\rho_{0}}{2D_{\mathrm{r}}}\alpha n_{0}(\alpha n_{0}-\rho_{0}\zeta)$ (A28)